# Preprint submitted to Journal

### v4

@@ 10.9 Optimization of Cartesian Product
However, currently, if we forcefully traverse the results of the Cartesian product, it is indeed possible to encounter the problem of computational explosion. Therefore, the optimization of Cartesian product calculation is also a research-worthy issue.
@@ Appendix A The simulation of $\mathrm{OIE}$ and 2 sequential operations
The simulated code implementation of complete sequential addition(Definition 20):
https://github.com/yleen/OIE_simulation/blob/master/src/simulation/OP/addition/complete_sequential_addition.py
The simulated code implementation of complete sequential multiplication(Definition 22):
https://github.com/yleen/OIE_simulation/blob/master/src/simulation/OP/multiplication/complete_sequential_multiplication.py
@@
https://github.com/yleen/OIE_simulation/blob/master/src/main.py

### v5

@@ 9.2 Comparison with other physical or philosophical understandings of simultaneity
Wang,Shouren “When you look at flowers, they bloom; when you don’t look, they fade into nothingness.” Simultaneity and event existence depend on conscious perception—events gain determinacy only through the observer’s awareness. $\mathrm{OIE}$ rejects pure perceptual dependence but retains the “subjective constructive” core.
@@ 10.9 Optimization of Cartesian Product
However, if someone forcefully needs to traverse the results of the Cartesian product, numerous existing optimization schemes are available for Cartesian product computation, such as the application of database technologies and various pruning algorithms. While we do not recommend allocating computing resources to Cartesian product calculations when using the Optional Intervals Event framework, it remains a worthwhile research topic to explore how to introduce database-based solutions, or alternatively, integrate the OIE framework and sequential operations into database technologies, if Cartesian product operations are deemed necessary for sequential computations in a handful of specific scenarios.
@@ Appendix A The simulation of $\mathrm{OIE}$ and 2 sequential operations
(Property 3) Ordering of starting timestamp and ending timestamp:
(Definition 4) Event with undetermined interval:
(Definition 6) Optional Intervals Event:
(Definition 7) The bound 2-tuple of a non-empty finite $\mathrm{2Tuple}\mathcal{T}$ instance:
(Property 5) The relationship between the $\mathcal{F}$(2nd element) and $\mathcal{I}$(3rd element) of $\mathrm{OIE}$:
(Definition 8) Equality of OIE instances:
(Definition 11) Set of $\mathcal{I}$ of finite $\mathrm{OIE}\mathcal{S}$ instance:
(Definition 12) Void $\mathrm{OIE}$:
(Definition 14)Function to get the set that is naturally isomorphic to the Cartesian product of all members of a finite $\mathrm{2Tuple}\mathcal{SS}$ instance in an index order:
(Definition 15) Set of tuples of infeasible interval 2-tuples of all members of a finite $\mathrm{OIE}\mathcal{S}$ instance under an index order:
(Definition 18 / function 1) Get the minimum 1st item of a $\mathrm{2Tuple}\mathcal{T}$ instance:
(Definition 18 / function 2) Get the maximum 2nd item of a $\mathrm{2Tuple}\mathcal{T}$ instance:
(Definition 19) Domain-filtered subset of a non-empty set of tuples of 2-tuples under a domain-filtering 2-tuple:
(Definition 20) Complete Sequential Addition:
(Definition 21) Ascending ordered filtered subset of the set of tuples of 2-tuples:
(Definition 22) Complete Sequential Multiplication:
(Property 11) Sequential operations involving identical $\mathrm{OIE}$ instances result in an $\mathrm{OIE}_{\mathrm{void}}$ instance:
(Property 12) The sequential operations involving $\mathrm{OIE}_{\mathrm{void}}$ instances result $\mathrm{oie}_{\mathrm{void}}$:
(Definition 24) Natural Complete Sequential Addition:
(Definition 27) The 1st type of implementation of an OIE instance:
(Definition 28) The 2nd type of implementation of an OIE instance: